# Multiobjective Aerodynamic Design Optimization of the NASA Common Research Model

### v1

@@ Abstract
Aircraft aerodynamic design optimization must account for the varying operating conditions along the cruise segment as opposed to designing at one fixed operating condition, to arrive at more realistic designs. Conventional approaches address this by performing a “multi-point” optimization that assumes a weighted average of the objectives at a set of sub-segments along the cruise segment. We argue that since such multi-point approaches are, inevitably, biased by the specification of the weights, they can lead to sub-optimal designs. Instead, we propose to optimize the aircraft design at multiple sub-segments simultaneously – that is, via multiobjective optimization that leads to a set of Pareto optimal solutions. However, existing work in multiobjective optimization suffers from (i) lack of sample efficiency (that is, keeping the number of function evaluations to convergence minimal), (ii) scalability with input dimensions and number of objectives, and (iii) the ability to generate a batch of iterates for synchronous parallel evaluations. To overcome these limitations, we propose a novel multiobjective Bayesian optimization methodology that demonstrates improved sample efficiency and accuracy compared to the state of the art. Inspired by Thompson sampling, our approach leverages Gaussian process surrogates and Bayesian decision theory to generate a sequence of iterates according to the probability that they are Pareto optimal. Our approach, named batch Pareto optimal Thompson sampling ($q\texttt{POTS}$), demonstrates superior empirical performance on a variety of synthetic experiments as well as a $24$ dimensional two-objective aerodynamic design optimization of the NASA common research model. We also provide open-source software of our methodology.
@@ 1 Introduction
Aerodynamic design optimization refers to optimizing the wing and control surface cross-sectional shapes, typically to minimize drag or maximize the lift-to-drag ratio (aka, the aerodynamic efficiency). As a result, it forms a routine part of the detailed design phase of aircraft design when the major design decisions are locked in. Conventionally, the aircraft wing is optimized for the cruise condition, which is reasonable given that it is the most dominant segment of a commercial aircraft’s mission. In this regard, it is quite common to consider a constant operating condition for the cruise segment. In principle, this is not a valid assumption because the aircraft weight is constantly decreasing along this segment (due to fuel consumption), resulting in changing required thrust, which in turn changes the operating angle of attack, Mach and Reynolds numbers; we provide a graphical illustration in Figure 1.
@@
Figure 1: A simplified mission profile for a commercial aircraft that emphasizes the cruise segment. The change in aircraft weight during the cruise segment induces change in required lift and thrust, and hence the operating angle of attack, Mach and Reynolds numbers. We anticipate that designs optimized for any one point along the cruise segment will likely be sub-optimal for other points. Therefore, we argue that “mulitobjective” optimization, that simultaneously considers multiple points along the cruise segment is necessary.
@@
In principle, one must consider a range of angles of attack, Mach, and Reynolds numbers in optimizing the wing shape. The aerodynamic design optimization discussion group (ADODG)  [^1] https://sites.google.com/view/mcgill-computational-aerogroup/adodg, accordingly, propounds $3$ and $9$ point “multi-point“ design optimization, which considers a weighted average over a set of operating conditions from the cruise segment. Those problems, however, involve a predefined weight for each sub-segment, which would need some domain knowledge to be specified appropriately. In this work, we argue that each sub-segment for the cruise must be simultaneously optimized, as opposed to scalarizing them using a bunch of weights. This is for the following reasons. First, we anticipate that the designs at individual sub-segments are conflicting with each other – that is, the optimal design at one segment can be sub-optimal at another. Second, the choice of weights, inevitably, affects the final design and hence must be carefully chosen, which might not be practical. Finally, simultaneously optimizing all objectives recovers an ensemble of solutions, of which the scalarized “multi-point” solution is one. We illustrate this on a synthetic test function, the ZDT3 test function, in Figure 2. Such a simultaneous optimization of a set of conflicting objectives is called multiobjective optimization, which forms the focus of this work. We provide a formal introduction in Section 2.
@@
Multiobjective optimization has been widely adopted for aerodynamic design optimization in the past. Evolutionary algorithms, particularly genetic algorithms, are arguably the most popular. For instance, Vicini and Quagliarella [55] present a numerical procedure for transonic airfoil design using genetic algorithms with single-point, multi-point, and multiobjective optimization formulations. The study highlights the advantages of multiobjective optimization over single-point and multi-point approaches in optimizing drag rise characteristics of transonic airfoils. Similarly, Coello Coello and Christiansen [9] introduce two new multiobjective optimization techniques based on genetic algorithms, demonstrating their ability to generate better trade-offs compared to other approaches. Advanced techniques such as scatter search and adaptive range multiobjective genetic algorithms have been applied to complex aerodynamic design problems. Molina et al. [35] describe a metaheuristic procedure based on scatter search for approximating the efficient frontier of nonlinear multiobjective optimization problems, showing its viability as an alternative for multiobjective optimization. Chiba et al. [8] report a large-scale application of evolutionary multiobjective optimization to the multidisciplinary design optimization of a transonic regional-jet aircraft wing, employing high-fidelity evaluation models and an adaptive range multiobjective genetic algorithm.
@@
Another angle is hybrid approaches combining evolutionary algorithms with surrogate models that can offer a sample-efficient (that is, minimal number of evaluations of the objectives) approach to multiobjective optimization. Leifsson et al. [33] present a computationally efficient procedure for multiobjective design optimization using variable-fidelity models and response surface surrogates. This approach combines an evolutionary algorithm with Kriging interpolation and space-mapping correction, allowing for efficient exploration-exploitation of trade-offs between conflicting objectives in airfoil design. Kim et al. [30] also demonstrate the effectiveness of a hybrid multiobjective evolutionary algorithm coupled with a response surface approximation surrogate model for axial-flow ventilation fan design, achieving significant improvements in efficiency and pressure rise. Keane [29] presents an approach that enables the parallel evolution of multiobjective Pareto sets, providing a statistically coherent means of solving expensive multiobjective design problems.
@@
Despite the advances made in the past, multiobjective aerodynamic design optimization still requires further attention. First, existing algorithms, such as evolutionary approaches, can incur a large number of objective evaluations, which can be prohibitive when using expensive high-fidelity aerodynamic models. Second, the efficient algorithms that exist today require derivative information, which can be difficult or impossible to obtain in several practical scenarios; see [39] for a detailed discussion. Finally, existing work scales poorly with the input dimensionality and the number of objectives. This work aims to address these limitations. Specifically, we take a surrogate-based approach, where we fit global surrogate models for our objectives. Then, via Bayesian decision theory, we adaptively enrich the surrogate model until a suitable convergence criterion is satisfied – this, in turn, results in a desirable approximation to the solution to the multiobjective optimization problem. Overall, our approach falls under the Bayesian optimization (BO) [20, 47] family of methods, and we provide more detail in Section 3. BO has been successfully demonstrated in the past to be competitive with the classical adjoint-based methods for aerodynamic design optimization [44, 39].
@@
The rest of the paper is organized as follows. In Section 2 we formally introduce our multiobjective optimization problem of interest. In Section 3 we present our overall methodology, and in Section 4 we provide details on our experiments and discuss results. Finally, we conclude in Section 5 with an outlook for future work.
@@ 2 Problem statement and preliminaries
Mathematically, we consider the simultaneous constrained optimization of $K$ objectives
$\begin{split}\max_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}~{}&\{f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})\}\end{split}$ (1)
@@
where $\mathbf{x}\in\boldsymbol{\mathcal{X}}\subset\mathbb{R}^{d}$ is the design variable, $f_{k}:\boldsymbol{\mathcal{X}}\rightarrow\mathbb{R},~{}\forall k=1,\ldots,K$, are expensive zeroth-order oracles (i.e., no available derivative information). $\boldsymbol{\mathcal{X}}$ is the domain of $f$. We are interested in identifying a Pareto optimal set of solutions, which “Pareto dominates” all other points. Let $\boldsymbol{f(\mathbf{x})}=[f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})]^{\top}$; when a solution $\boldsymbol{f(\mathbf{x})}$ Pareto dominates another solution $\boldsymbol{f(\mathbf{x}^{\prime})}$, then $f_{k}(\mathbf{x})\geq f_{k}(\mathbf{x}^{\prime}),~{}\forall k=1,\ldots,K$ and $\exists k\in[K]$ such that $f_{k}(\mathbf{x})>f_{k}(\mathbf{x}^{\prime})$. We write Pareto dominance as $\boldsymbol{f(\mathbf{x})}\succ\boldsymbol{f(\mathbf{x}^{\prime})}$. The set $\mathcal{Y}^{*}=\{\boldsymbol{f(\mathbf{x})}~{}:~{}\nexists\mathbf{x}^{\prime}\in\boldsymbol{\mathcal{X}}:\boldsymbol{f(\mathbf{x}^{\prime})\succ\boldsymbol{f(\mathbf{x})}}\}$ is called the Pareto frontier and the set $\boldsymbol{\mathcal{X}}^{*}=\{\mathbf{x}\in\boldsymbol{\mathcal{X}}~{}:~{}\boldsymbol{f(\mathbf{x})}\in\mathcal{Y}^{*}\}$ is called the Pareto set.
@@
We focus on sample-efficient methods to address our problem. Evolutionary algorithms (EAs), such as the nondominated sorting genetic algorithm (NSGA)[16, 17], are widely used for multiobjective optimization (see [65] for a review). However, EA-based methods typically require a large number of evaluations, which can be impractical in settings involving expensive function evaluations. To overcome this limitation, a common strategy is to employ Bayesian optimization (BO) with Gaussian process (GP) surrogate models for each objective. In this approach, a GP is trained for each objective using an initial dataset $\boldsymbol{\mathcal{D}}^{i}_{n}={(\mathbf{x}_{j},y_{j}),j=1,\ldots,n}$, where $y_{j}=f_{i}(\mathbf{x}_{j})$. An acquisition function is then constructed to measure the utility of a candidate point $\mathbf{x}$, and a subproblem is solved to identify the point(s) that maximize this acquisition function. The surrogate models are subsequently updated with observations at the new points, and this process is iterated until a chosen convergence condition is satisfied. We formally introduce this methodology in Section 2.1. BO is especially popular in the single-objective case ($K=1$); for example, the expected improvement (EI)[28] and probability of improvement (PI)[34, 27] quantify the likelihood that a candidate point improves upon the best observed value. The GP upper confidence bound (UCB)[48] offers a conservative estimate of the maximum by leveraging the GP’s mean and uncertainty estimates. Other notable acquisition strategies for $K=1$ include entropy-based methods[56, 24], the knowledge gradient (KG)[21, 58, 59], and stepwise uncertainty reduction (SUR)[37, 7]. Multiobjective Bayesian optimization (MOBO) generalizes BO to handle multiple objectives; we review relevant prior work in the sections that follow.
@@ 2.1 Single objective Bayesian optimization
We place a Gaussian process (GP) prior on the oracle, $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\cdot))$, where $k(\cdot,\cdot):\boldsymbol{\mathcal{X}}\times\boldsymbol{\mathcal{X}}\rightarrow\mathbb{R}+$ denotes the covariance function (also known as the kernel). Observations from the oracle are modeled as $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ for $i=1,\ldots,n$, where each $\epsilon_{i}$ is assumed to be zero-mean Gaussian noise with unknown variance $\tau^{2}$. Using the observed dataset $\boldsymbol{\mathcal{D}}_{n}=\{\mathbf{x}_{i},y_{i},\tau^{2}\}_{i=1}^{n}$, we fit a posterior GP—parameterized by hyperparameters $\boldsymbol{\Omega}$—to obtain the following conditional posterior distribution [42]:
@@
An alternative to constructing acquisition functions is Thompson sampling (TS)[51], which selects actions based on the probability that they are optimal in a specified sense[45]. In the context of single-objective optimization, TS selects points according to the distribution $p_{\mathbf{x}^{*}}(\mathbf{x})$, where $\mathbf{x}^{*}$ denotes the maximizer of the objective function. Within BO, this distribution can be written as:
[…]

### v2

@@ Abstract
Aircraft aerodynamic design optimization must account for the varying operating conditions along the cruise segment as opposed to designing at one fixed operating condition, to arrive at more realistic designs. Conventional approaches address this by performing a “multi-point” optimization that assumes a weighted average of the objectives at a set of sub-segments along the cruise segment. We argue that since such multi-point approaches are, inevitably, biased by the specification of the weights, they can lead to sub-optimal designs. Instead, we propose to optimize the aircraft design at multiple sub-segments simultaneously – that is, via multiobjective optimization that leads to a set of Pareto optimal solutions. However, existing work in multiobjective optimization suffers from (i) lack of sample efficiency (that is, keeping the number of function evaluations to convergence minimal), (ii) scalability in the absence of derivative information, and (iii) the ability to generate a batch of iterates for synchronous parallel evaluations. To overcome these limitations, we apply a novel multiobjective Bayesian optimization methodology for aerodynamic design optimization that demonstrates improved sample efficiency and accuracy compared to the state of the art. Inspired by Thompson sampling, our approach leverages Gaussian process surrogates and Bayesian decision theory to generate a sequence of iterates according to the probability that they are Pareto optimal. Our approach, named batch Pareto optimal Thompson sampling ($q\texttt{POTS}$) [^1] Here, $q$ stands for selecting a batch of $q$ iterates at every step., demonstrates superior empirical performance on a variety of synthetic experiments as well as a $24$ dimensional two-objective aerodynamic design optimization of the NASA common research model. We also provide open-source software of our methodology and experiments.
keywords:
Aerodynamic design optimization , multiobjective Bayesian optimization, Gaussian process regression
@@ 1 Introduction
Aerodynamic design optimization refers to optimizing the wing and control surface cross-sectional shapes, typically to minimize drag or maximize the lift-to-drag ratio (aka, the aerodynamic efficiency). As a result, it forms a routine part of the detailed design phase of aircraft design when the major design decisions are locked in. Conventionally, the aircraft wing is optimized for the cruise condition, which is reasonable given that it is the most dominant segment of a commercial aircraft’s mission. In this regard, it is quite common to consider a constant operating condition for the cruise segment. In principle, this is not a valid assumption because the aircraft weight is constantly decreasing along this segment (due to fuel consumption), resulting in changing required thrust, which in turn changes the operating angle of attack, Mach and Reynolds numbers; we provide a graphical illustration in Figure˜1.
@@
Figure 1: A simplified mission profile for a commercial aircraft that emphasizes the cruise segment. The change in aircraft weight during the cruise segment induces change in required lift and thrust, and hence the operating angle of attack, Mach and Reynolds numbers. We anticipate that designs optimized for any one point along the cruise segment will likely be sub-optimal for other points. Therefore, we argue that “multiobjective” optimization, that simultaneously considers multiple points along the cruise segment is necessary.
@@
In principle, one must consider a range of angles of attack, Mach, and Reynolds numbers in optimizing the wing shape. The aerodynamic design optimization discussion group (ADODG)  [^2] https://sites.google.com/view/mcgill-computational-aerogroup/adodg, accordingly, propounds $3$ and $9$ point “multi-point“ design optimization, which considers a weighted average over a set of operating conditions from the cruise segment. Those problems, however, involve a predefined weight for each sub-segment, which would need some domain knowledge to be specified appropriately. In this work, we argue that each sub-segment for the cruise segment must be simultaneously optimized, as opposed to scalarizing them using a bunch of weights. This is for the following reasons. First, we anticipate that the designs at individual sub-segments are conflicting with each other – that is, the optimal design at one segment can be sub-optimal at another. Second, the choice of weights, inevitably, affects the final design and hence must be carefully chosen, which might not be practical. Finally, simultaneously optimizing all objectives recovers an ensemble of solutions, of which the scalarized “multi-point” solution is one. Specifically, minimizing a weighted sum with strictly positive weights always yields a Pareto–optimal solution (hence a scalarized solution is on the Pareto front). However, the converse requires convexity of the attainable objective set: in nonconvex problems linear scalarizations cannot recover Pareto points in concave regions of the front, which motivates the use of direct multiobjective methods like ours. We illustrate this on a synthetic test function, the ZDT3 test function, in Figure˜2, and provide a concrete theoretical justification in Section˜5.2.
@@
Such a simultaneous optimization of a set of conflicting objectives is called multiobjective optimization, which forms the focus of this work – we provide a formal introduction in Section˜2. This is crucial for aircraft design because the design space can include regions of both conflicting and nonconflicting designs; this is shown in Figure˜3, where we consider the aircraft drag at two different cruise points, which are to be minimized. When designs are nonconflicting (as in the bottom right of Figure˜3), then a multipoint optimum, regardless of the choice of weights, will find the same optimal design. However, when the designs are conflicting, a “Pareto optimal” set of designs must be recovered (red circles in Figure˜3). Therefore, efficient methods to simultaneously optimize objectives at multiple operating conditions are crucial – this work makes contributions in that direction.
Figure 3: Plot shows aircraft drag for the NASA CRM evaluated at two different points along cruise segment for a variety of wing shapes. Notice that certain designs are nonconflicting, while others are not. Multiobjective optimization is necessary to properly account for design trade-offs and identify a set of optimal solutions for a practitioner to choose from.
@@
Multiobjective optimization has been widely adopted for aerodynamic design optimization in the past. Evolutionary algorithms, particularly genetic algorithms, are arguably the most popular. For instance, vicini_inverse_1997 present a numerical procedure for transonic airfoil design using genetic algorithms with single-point, multi-point, and multiobjective optimization formulations. The study highlights the advantages of multiobjective optimization over single-point and multi-point approaches in optimizing drag rise characteristics of transonic airfoils. Similarly, coello_coello_moses:_1999 introduce two new multiobjective optimization techniques based on genetic algorithms, demonstrating their ability to generate better trade-offs compared to other approaches. Advanced techniques such as scatter search and adaptive range multiobjective genetic algorithms have been applied to complex aerodynamic design problems. molina_sspmo:_2007 describe a metaheuristic procedure based on scatter search for approximating the efficient frontier of nonlinear multiobjective optimization problems, showing its viability as an alternative for multiobjective optimization. chiba_multidisciplinary_2007 report a large-scale application of evolutionary multiobjective optimization to the multidisciplinary design optimization of a transonic regional-jet aircraft wing, employing high-fidelity evaluation models and an adaptive range multiobjective genetic algorithm.
@@
Another angle is hybrid approaches combining evolutionary algorithms with surrogate models that can offer a sample-efficient (that is, minimal number of evaluations of the objectives) approach to multiobjective optimization. leifsson_multiobjective_2015 present a computationally efficient procedure for multiobjective design optimization using variable-fidelity models and response surface surrogates. This approach combines an evolutionary algorithm with Kriging interpolation and space-mapping correction, allowing for efficient exploration-exploitation of trade-offs between conflicting objectives in airfoil design. kim_axial-flow_2011 also demonstrates the effectiveness of a hybrid multiobjective evolutionary algorithm coupled with a response surface approximation surrogate model for axial-flow ventilation fan design, achieving significant improvements in efficiency and pressure rise. keane_statistical_2006 presents an approach that enables the parallel evolution of multiobjective Pareto sets, providing a statistically coherent means of solving expensive multiobjective design problems. Other recent efforts include [zuhal2019comparative] and [lin2023parallel].
@@
Despite the advances made in the past, multiobjective aerodynamic design optimization still requires further attention. First, existing algorithms, such as evolutionary approaches, can incur a large number of objective evaluations, which can be prohibitive when using expensive high-fidelity aerodynamic models. Second, the efficient algorithms that exist today require derivative information, which can be difficult or impossible to obtain in several practical scenarios; see [plaban2025derivative] for a detailed discussion. Finally, in existing work sampling a batch of iterates at every step, for asynchronous parallel evaluations on high-performance computing environments, is either not possible or incurs a nontrivial additional cost. This work aims to address these limitations. Specifically, we take a surrogate-based approach, where we fit global surrogate models for our objectives. Then, via Bayesian decision theory, we adaptively enrich the surrogate model until a suitable convergence criterion is satisfied – this, in turn, results in a desirable approximation to the solution to the multiobjective optimization problem. Overall, our approach falls under the Bayesian optimization (BO) [frazier2018tutorial, shahriari2015taking] family of methods, and we provide more detail in Section˜3. BO has been successfully demonstrated in the past to be competitive with the classical adjoint-based methods for aerodynamic design optimization [renganathan2021enhanced, plaban2025derivative].
In this work, we argue that multiobjective aerodynamic design optimization is more effective than solving multiple multi-point optimization problems. This is because, a single solution of a multiobjective optimization problem results in a set of designs that covers every possible multi-point design. The converse, however, is not true: multi-point optimization, regardless of the scalarizing weights, is not guaranteed to find a Pareto optimal point. We substantiate this argument theoretically in Section˜5.2. Therefore, naturally, direct multiobjective optimization methods are necessary. In this regard, to address the existing limitations in solving multiobjective aerodynamic design optimization problems in the absence of adjoint information, we propose a novel multiobjective Bayesian optimization method and demonstrate it on the aerodynamic design optimization of the NASA common research model (CRM).
Remark 1. As a first step, our study is scoped down to unconstrained multiobjective aerodynamic design optimization of the NASA CRM under inviscid flow conditions; we briefly discuss constrained extensions in Section˜5.1. However, we rigorously benchmark the proposed unconstrained optimization method against the state of the art to establish the merits of the method.
Remark 2. In the current work, we explore multiobjective optimization of modest input (2242-24) and output dimensionalities ($2$); high-dimensional scaling is reserved for future work. Demonstration on fully viscous constrained multiobjective aerodynamic design is reserved for future work.
@@
The rest of the paper is organized as follows. In Section˜2 we formally introduce our multiobjective optimization problem of interest. In Section˜3 we present our overall methodology, and in Section˜4 we provide details on our experiments and discuss results. Finally, we conclude in Section˜5 with an outlook for future work.
@@ 2 Problem statement and preliminaries
Mathematically, we consider the simultaneous unconstrained optimization of $K$ objectives
$\begin{split}\max_{\mathbf{x}\in\mathcal{X}}~&\{f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})\}\end{split}$ (1)
@@
where $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$ is the design variable, $f_{k}:\mathcal{X}\rightarrow\mathbb{R},~\forall k=1,\ldots,K$, are expensive zeroth-order oracles (i.e., no available derivative information). $\mathcal{X}$ is the domain of $f$. We are interested in identifying a Pareto optimal set of solutions, which “Pareto dominates” all other points. Let $\boldsymbol{f(\mathbf{x})}=[f_{1}(\mathbf{x}),\ldots,f_{K}(\mathbf{x})]^{\top}$; when a solution $\boldsymbol{f(\mathbf{x})}$ Pareto dominates another solution $\boldsymbol{f(\mathbf{x}^{\prime})}$, then $f_{k}(\mathbf{x})\geq f_{k}(\mathbf{x}^{\prime}),~\forall k=1,\ldots,K$ and $\exists k\in[K]$ such that $f_{k}(\mathbf{x})>f_{k}(\mathbf{x}^{\prime})$. We write Pareto dominance as $\boldsymbol{f(\mathbf{x})}\succ\boldsymbol{f(\mathbf{x}^{\prime})}$. The set $\mathcal{Y}^{*}=\{\boldsymbol{f(\mathbf{x})}~:~\nexists\mathbf{x}^{\prime}\in\mathcal{X}:\boldsymbol{f(\mathbf{x}^{\prime})\succ\boldsymbol{f(\mathbf{x})}}\}$ is called the Pareto frontier and the set $\mathcal{X}^{*}=\{\mathbf{x}\in\mathcal{X}~:~\boldsymbol{f(\mathbf{x})}\in\mathcal{Y}^{*}\}$ is called the Pareto set.
@@
We focus on sample-efficient methods to address our problem. Evolutionary algorithms (EAs), such as the nondominated sorting genetic algorithm (NSGA)[deb2000fast, deb2002fast], are widely used for multiobjective optimization (see [zitzler2000comparison] for a review). However, EA-based methods typically require a large number of evaluations, which can be impractical in settings involving expensive function evaluations. To overcome this limitation, a common strategy is to employ Bayesian optimization (BO) with Gaussian process (GP) surrogate models for each objective. In this approach, a GP is trained for each objective using an initial dataset $\boldsymbol{\mathcal{D}}^{i}_{n}={(\mathbf{x}_{j},y_{j}),j=1,\ldots,n}$, where $y_{j}=f_{i}(\mathbf{x}_{j})$. An acquisition function is then constructed to measure the utility of a candidate point $\mathbf{x}$, and a subproblem is solved to identify the point(s) that maximize this acquisition function. The surrogate models are subsequently updated with observations at the new points, and this process is iterated until a chosen convergence condition is satisfied. We formally introduce this methodology in Section˜2.1. BO is especially popular in the single-objective case ($K=1$); for example, the expected improvement (EI)[jones1998efficient] and probability of improvement (PI)[mockus1978application, jones2001taxonomy] quantify the likelihood that a candidate point improves upon the best observed value. The GP upper confidence bound (UCB)[srinivas2009gaussian] offers a conservative estimate of the maximum by leveraging the GP’s mean and uncertainty estimates. Other notable acquisition strategies for $K=1$ include entropy-based methods[wang2017max, hernandez2014predictive], the knowledge gradient (KG)[frazier2008knowledge, wu2016parallel, wu2019practical], and stepwise uncertainty reduction (SUR)[picheny2014stepwise, chevalier2014fast]. Multiobjective Bayesian optimization (MOBO) generalizes BO to handle multiple objectives; we review relevant prior work in the sections that follow.
@@ 2.1 Single objective Bayesian optimization
We place a Gaussian process (GP) prior on the oracle, $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\cdot))$, where $k(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}+$ denotes the covariance function (also known as the kernel). Observations from the oracle are modeled as $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ for $i=1,\ldots,n$, where each $\epsilon_{i}$ is assumed to be zero-mean Gaussian noise with unknown variance $\tau^{2}$. Using the observed dataset $\boldsymbol{\mathcal{D}}_{n}=\{\mathbf{x}_{i},y_{i},\tau^{2}\}_{i=1}^{n}$, we fit a posterior GP—parameterized by hyperparameters $\boldsymbol{\Omega}$—to obtain the following conditional posterior distribution [rasmussen:williams:2006]:
@@
An alternative to constructing acquisition functions is Thompson sampling (TS)[thompson1933likelihood], which selects actions based on the probability that they are optimal in a specified sense[russo2014learning]. In the context of single-objective optimization, TS selects points according to the distribution $p_{\mathbf{x}^{*}}(\mathbf{x})$, where $\mathbf{x}^{*}$ denotes the maximizer of the objective function. Within BO, this distribution can be written as:
$\begin{split}p_{\mathbf{x}^{*}}(\mathbf{x})=&\int p_{\mathbf{x}^{*}}(\mathbf{x}\mid Y)p(Y\mid\boldsymbol{\mathcal{D}}_{n})dY=\int\delta\left(\mathbf{x}-\arg\max_{\mathbf{x}\in\mathcal{X}}Y(\mathbf{x})\right)p(Y\mid\boldsymbol{\mathcal{D}}_{n})dY,\end{split}$ (3)